# Meta-Homogenization for Knitwear Simulation

### v2

@@ 6.2. Adjoint Gauss-Newton
(19) $\bm{H}_{i}=\frac{\partial^{2}\epsilon}{\partial\bm{\gamma}^{2}}=\left(\frac{\partial\bm{x}^{M}_{i}}{\partial\bm{\gamma}}\right)^{\top}\cdot\frac{\partial^{2}\epsilon_{i}}{\partial\bm{x}^{M^{2}}_{i}}\cdot\frac{\partial\bm{x}^{M}_{i}}{\partial\bm{\gamma}}+\frac{\partial\epsilon_{i}}{\partial\bm{x}^{M}_{i}}\cdot\frac{\partial^{2}\bm{x}_{i}^{M}}{\partial\bm{\gamma}^{2}},$
@@
We observe that when $\epsilon_{i}$ is small and mesh deformation $\bm{x}^{M}_{i}$ starts aligning with the target yarn pose $\bm{x}^{Y}$, $\partial\epsilon_{i}/\partial\bm{x}^{M}_{i}$ is always close to zero. This allows an alternative precondition matrix $\bm{P}_{i}$ that discards the second term of $\mathbf{H}_{i}$, such that
(20) $\bm{P}_{i}=\left(\frac{\partial\bm{x}^{M}_{i}}{\partial\bm{\gamma}}\right)^{\top}\bm{G}_{i}\frac{\partial\bm{x}^{M}_{i}}{\partial\bm{\gamma}}\;,$
where, $\bm{G}_{i}={\partial^{2}\epsilon_{i}}/{\partial\bm{x}^{M^{2}}_{i}}\in\mathbb{R}^{3n^{M}\times 3n^{M}}$ and can be pre-computed. This Hessian simplification strategy is a.k.a. Gauss-Newton method, and the preconditioned search direction can be obtained by solving the linear system $\bm{P}_{i}$:
@@
While the formulation is well known, Gauss-Newton is seldom used with an adjoint method in practice. The difficulty lies in the fact that $\partial\bm{x}^{M}_{i}/\partial\bm{\gamma}$ should never be explicitly evaluated via solving Eq. 16 and, therefore, we do not have the actual precondition matrix $\bm{P}_{i}$. Even if we could compute $\bm{P}_{i}$ exactly, solving Eq. 21 remains prohibitive since $\bm{P}_{i}$ is likely a dense matrix.
We propose a novel formulation to tackle this challenge. Specifically, we introduce two more adjoint state vectors $\bm{\mu},\bm{\nu}\in\mathbb{R}^{3n^{M}}$ and form the following linear system:
@@
(22) $\left[\begin{array}[]{ccc}\bm{G}_{i}&-\left(\frac{\partial\bm{g}}{\partial\bm{x}^{M}_{i}}\right)^{\top}&\bm{0}\\ -\frac{\partial\bm{g}}{\partial\bm{x}^{M}_{i}}&\bm{0}&\frac{\partial\bm{g}}{\partial\bm{\gamma}}\\ \bm{0}&\left(\frac{\partial\bm{g}}{\partial\bm{\gamma}}\right)^{\top}&\bm{0}\end{array}\right]\left[\begin{array}[]{c}\bm{\mu}\\ \bm{\nu}\\ \bm{d}_{i}^{GN}\end{array}\right]=\left[\begin{array}[]{c}\bm{0}\\ \bm{0}\\ \left(\frac{\partial\epsilon_{i}}{\partial\bm{\gamma}}\right)^{\top}\end{array}\right].$
@@
One can quickly verify that the solution of Eq. 22 coincides with the solution of Eq. 21: from the first two lines of Eq. 22, we have $(\partial\bm{g}/\partial\bm{x}^{M}_{i})^{-\top}\bm{G}_{i}\bm{\mu}=\bm{\nu}$ and $(\partial\bm{g}/\partial\bm{x}^{M}_{i})^{-1}(\partial\bm{g}/\partial\bm{\gamma})\bm{d}_{i}^{GN}=\bm{\mu}$, suggesting
@@
$\left(\frac{\partial\bm{g}}{\partial\bm{x}_{i}^{M}}\right)^{-\top}\bm{G}_{i}\underbrace{\left(\frac{\partial\bm{g}}{\partial\bm{x}_{i}^{M}}\right)^{-1}\frac{\partial\bm{g}}{\partial\bm{\gamma}}}_{-\partial\bm{x}_{i}^{M}/\partial\bm{\gamma}}\bm{d}_{i}^{GN}=\bm{\nu}.$
@@
Left multiplying $(\partial\bm{g}/\partial\bm{\gamma})^{\top}$ both sides and knowing $(\partial\bm{g}/\partial\bm{\gamma})^{\top}\bm{d}_{i}^{GN}=\partial\epsilon_{i}/\partial\bm{x}_{i}^{M}$, i.e. the third row of Eq. 22, restores the formula back to the vanilla Gauss-Newton of Eq. 21.
@@
When $\bm{\gamma}$ is fixed, $\epsilon_{i}$ becomes a quadratic form of $\bm{x}^{M}_{i}$, and $\bm{G}_{i}$ is therefore non-negative definite. To this end, we follow the Levenberg-Marquardt method (Moré, 2006) that adds a small diagonal $\kappa\bm{I}$ at l.h.s. of Eq. 22 to secure its positive definiteness. It is noteworthy that $\partial\bm{g}/\partial\bm{x}_{i}^{M}$ is a sparse matrix–it has non-zero entities only at the adjacent mesh elements (similar to FEM matrices), and $\partial\bm{g}/\partial\bm{\gamma}$ is also a sparse matrix (only the element’s material parameter influences its residual force). Solving Eq. 22 is more efficient than solving Eq. 21 despite the increased matrix dimension.
@@ 6.3. Linear Search & Non-Negativity Constraint
It should be noted that all the components i.e., $\partial\bm{g}/\partial\bm{x}^{M}_{i}$, $\partial\bm{g}/\partial\bm{\gamma}$, and $\partial\epsilon_{i}/\partial\bm{\gamma}$ that assemble the Gauss-Newton system are readily available after the gradient computation. Compared with gradient descent, the additional cost of our adjoint Gauss-Newton comes from the linear solve of Eq. 22. Thanks to this novel formulation and the sparsity of the matrix, this step is not the bottleneck of the pipeline. Instead, the most expensive computation is always at finding $\bm{x}^{M}_{i}$ to satisfy the equilibrium condition i.e., Eq. 15. As to be discussed in Sec. 7, the constraint-based energy formulation (Eq. 8) allows the use of highly efficient parallel GPU procedures to accelerate this step, which is not only helpful for the forward simulation but also for the elasticity fitting.
@@ 6.4. Sample-by-Sample Fitting
(23) $\bm{\gamma}\leftarrow\frac{w}{w+w_{k}}\bm{\gamma}+\frac{w_{k}}{w+w_{k}}\bm{\gamma}_{k},.$
@@ 6.5. Harmonic Initialization
Figure 4. Harmonic initialization.  We design a progressive initialization strategy based on Harmonic bases of the volume mesh. By projecting the material vector into the Harmonic subspaces of different ranks, meta-homogenization always finds a reliable initial value for the two-stage elasticity fitting procedure.
@@
For instance, to switch the optimization target to $\bm{q}$ in Eq. 22, we right multiply $\mathsf{diag}(\bm{H})$ to $\partial\bm{g}/\partial\bm{\gamma}$ and $(\partial\epsilon_{i}/\partial\bm{\gamma})^{\top}$, and sandwich $\bm{G}_{i}$ with $\mathsf{diag}(\bm{H})$ as
@@ 7.1. Projective Dynamics
The global step is a standard linear solve in the form of
(28) $\bm{K}\bm{x}=\bm{b},$
where we have
@@
The global step stands as the most expensive computation along the pipeline as a pre-factorized $\bm{K}$ is a large dense matrix that is not GPU-friendly.
@@ 7.2. Domain Decomposition
To accelerate the global step computation, we exploit the component mode synthesis (CMS) (Craig Jr and Hale, 1988; Craig Jr, 1985) to decompose the system into multiple domains (or components) and build a linear subspace at each domain to analyze the dynamic responses of complex structures. CMS was originally designed for linear structural analysis, and its generalization to nonlinear simulation remains an open research problem. Fortunately, the unique modality of PD allows us to apply CMS just at the global stage solve, so we can partition the mesh into multiple domains without aligning the decomposition with the variation of the underlying knit patterns.
Unknown DOFs at each domain can be now grouped into internal DOFs and boundary DOFs. As the name suggests, the boundary DOFs interface with the neighbor domains while the internal DOFs are isolated by the boundary DOFs.
Figure 7. Comparison with HYLC (1$\times$1 rib).  We compare our method with HYLC and homogenize a square knitted fabric of the periodic 1$\times$1 rib pattern. The full-scale YLS results are shown at the bottom for reference. In general, both our method and HYLC yield plausible results, and both are visually similar to YLS results. Meta-homogenization uses a mesh of $38$K DOFs, while HYLC only has $9$K DOFs. Nevertheless, our method takes 85ms per frame and is 180$\sim$ 180$\times$ 180 × faster than HYLC.
@@
Due to the linearity of this problem, $\bm{\Psi}_{(d)}^{(i)}$ encodes all the possible internal responses induced by boundary stimuli. When the domain does not undertake any non-boundary loads, $\bm{\Psi}_{(d)}^{(i)}$ relates boundary DOFs and internal DOFs as
(34) $\bm{x}^{(i)}_{(d)}=\bm{\Psi}_{(d)}^{(i)}\bm{x}^{(b)}_{(d)}.$
They are a.k.a. boundary modes in CMS, and serve as subspace basis vectors for our global solve. For non-boundary responses, we compute a compact set of eigenvectors (e.g., 20) of $\bm{K}_{(d)}^{(ii)}$ corresponding to the smallest eigenvalues
@@
Figure 8. Comparison with HYLC (stockinette).  Given fabric with a periodic stockinette pattern, our method effectively captures spatially varying deformation leveraging material heterogeneity, particularly under significant stretch as highlighted. In the example, meta-homogenization uses a mesh of $42$K DOFs, while HYLC has $12$K DOFs. Our method is 350$\sim$ 350$\times$ 350 × faster than HYLC.
Figure 9. Comparison with HYLC (draping).  We drape the ribbing patch on a sphere. Our method better replicates local deformation in the middle than HYLC. Note that YLS simulation result for the same draping test is not used for elasticity fitting, which demonstrates the generalizability of our method.
Domain decomposition and CMS-like subspace construction techniques make our pipeline less sensitive to global matrix change e.g., due to material update or the presence of new constraints of collisions. After the reduced global solve is obtained, we use Eq. 37 to convert the generalized coordinate to the fullspace result. This result is then sent to a GPU-based full-space solver to make sure the global solution is accurate. In our implementation, we use the aggregated Jacobi method (A-Jacobi) proposed in (Lan et al., 2022). A-Jacobi combines two or three Jacobi iterations into one iteration to fully exploit the parallel capacity of modern GPUs. As a result, mesh-level simulation is efficient is nearly interactive in many examples reported in this paper.
Note that a fast and numerically stable mesh simulation is critical to the entire pipeline–it is not only helpful for simulating the meta-homogenized knitwear materials but also an indispensable module for elasticity fitting. Recall that our adjoint Gauss-Newton is only feasible when Eq. 15 is satisfied. This requirement suggests we need to find the equilibrium configuration of the mesh given the current material vector $\bm{\gamma}$. Whenever $\bm{\gamma}$ are updated during the elasticity fitting process, we have to re-assemble and factorize $\bm{x}^{M}_{i}$, which is the most expensive computation along the pipeline. Without a fast solver, elasticity fitting at such a high-dimension material space is infeasible.

### v3

@@ 6.2. Adjoint Gauss-Newton
(18) $\bm{H}_{i}=\frac{\partial^{2}\epsilon}{\partial\bm{\gamma}^{2}}=\left(\frac{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}{\partial\bm{\gamma}}\right)^{\top}\cdot\frac{\partial^{2}\epsilon_{i}}{\partial\bm{x}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}^{2}}_{i}}\cdot\frac{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}{\partial\bm{\gamma}}+\frac{\partial\epsilon_{i}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}\cdot\frac{\partial^{2}\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}{\partial\bm{\gamma}^{2}},$
@@
We observe that when $\epsilon_{i}$ is small and mesh deformation $\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}$ starts aligning with the target yarn pose $\bm{x}^{Y}$, $\partial\epsilon_{i}/\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}$ is always close to zero. This allows an alternative precondition matrix $\bm{P}_{i}$ that discards the second term of $\mathbf{H}_{i}$, such that
(19) $\bm{P}_{i}=\left(\frac{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}{\partial\bm{\gamma}}\right)^{\top}\bm{G}_{i}\frac{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}{\partial\bm{\gamma}}\;,$
where, $\bm{G}_{i}={\partial^{2}\epsilon_{i}}/{\partial\bm{x}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}^{2}}_{i}}\in\mathbb{R}^{3n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\times 3n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}$ and can be pre-computed. This Hessian simplification strategy is a.k.a. Gauss-Newton method, and the preconditioned search direction can be obtained by solving the linear system $\bm{P}_{i}$:
@@
While the formulation is well known, Gauss-Newton is seldom used with an adjoint method in practice. The difficulty lies in the fact that $\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}/\partial\bm{\gamma}$ should never be explicitly evaluated via solving Eq. 15 and, therefore, we do not have the actual precondition matrix $\bm{P}_{i}$. Even if we could compute $\bm{P}_{i}$ exactly, solving Eq. 20 remains prohibitive since $\bm{P}_{i}$ is likely a dense matrix.
Zehnder et al. (2021) demonstrate that this issue can be circumvented. Similar to how the adjoint method avoids explicit computation of $\frac{\partial x^{M}_{i}}{\partial\gamma}$, two additional adjoint state vectors are introduced to transform the dense system Eq. 20 into the following larger sparse linear system:
@@
(21) $\left[\begin{array}[]{ccc}\bm{G}_{i}&-\left(\frac{\partial\bm{g}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}\right)^{\top}&\bm{0}\\ -\frac{\partial\bm{g}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}&\bm{0}&\frac{\partial\bm{g}}{\partial\bm{\gamma}}\\ \bm{0}&\left(\frac{\partial\bm{g}}{\partial\bm{\gamma}}\right)^{\top}&\bm{0}\end{array}\right]\left[\begin{array}[]{c}\bm{\mu}\\ \bm{\nu}\\ \bm{d}_{i}^{GN}\end{array}\right]=\left[\begin{array}[]{c}\bm{0}\\ \bm{0}\\ \left(\frac{\partial\epsilon_{i}}{\partial\bm{\gamma}}\right)^{\top}\end{array}\right].$
@@
One can quickly verify that the solution of Eq. 21 coincides with the solution of Eq. 20: from the first two lines of Eq. 21, we have $(\partial\bm{g}/\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i})^{-\top}\bm{G}_{i}\bm{\mu}=\bm{\nu}$ and $(\partial\bm{g}/\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i})^{-1}(\partial\bm{g}/\partial\bm{\gamma})\bm{d}_{i}^{GN}=\bm{\mu}$, suggesting
@@
$\left(\frac{\partial\bm{g}}{\partial\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}\right)^{-\top}\bm{G}_{i}\underbrace{\left(\frac{\partial\bm{g}}{\partial\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}\right)^{-1}\frac{\partial\bm{g}}{\partial\bm{\gamma}}}_{-\partial\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}/\partial\bm{\gamma}}\bm{d}_{i}^{GN}=\bm{\nu}.$
@@
Left multiplying $(\partial\bm{g}/\partial\bm{\gamma})^{\top}$ both sides and knowing $(\partial\bm{g}/\partial\bm{\gamma})^{\top}\bm{\nu}=\partial\epsilon_{i}/\partial\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$, i.e. the third row of Eq. 21, restores the formula back to the vanilla Gauss-Newton of Eq. 20.
@@
When $\bm{\gamma}$ is fixed, $\epsilon_{i}$ becomes a quadratic form of $\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}$, and $\bm{G}_{i}$ is therefore non-negative definite. To this end, we follow the Levenberg-Marquardt method (Moré, 2006) that adds a small diagonal $\kappa\bm{I}$ at l.h.s. of Eq. 21 to secure its positive definiteness. It is noteworthy that $\partial\bm{g}/\partial\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ is a sparse matrix–it has non-zero entities only at the adjacent mesh elements (similar to FEM matrices), and $\partial\bm{g}/\partial\bm{\gamma}$ is also a sparse matrix (only the element’s material parameter influences its residual force). Solving Eq. 21 is more efficient than solving Eq. 20 despite the increased matrix dimension.
@@ 6.3. Linear Search & Non-Negativity Constraint
It should be noted that all the components i.e., $\partial\bm{g}/\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}$, $\partial\bm{g}/\partial\bm{\gamma}$, and $\partial\epsilon_{i}/\partial\bm{\gamma}$ that assemble the Gauss-Newton system are readily available after the gradient computation. Compared with gradient descent, the additional cost of our adjoint Gauss-Newton comes from the linear solve of Eq. 21. Thanks to the sparsity of the matrix, this step is not the bottleneck of the pipeline. Instead, the most expensive computation is always at finding $\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}$ to satisfy the equilibrium condition i.e., Eq. 14. As to be discussed in Sec. 7, the constraint-based energy formulation (Eq. 7) allows the use of highly efficient parallel GPU procedures to accelerate this step, which is not only helpful for the forward simulation but also for the elasticity fitting.
@@ 6.4. Sample-by-Sample Fitting
(22) $\bm{\gamma}\leftarrow\frac{w}{w+w_{k}}\bm{\gamma}+\frac{w_{k}}{w+w_{k}}\bm{\gamma}_{k},$
@@ 6.5. Harmonic Initialization
Figure 4. Harmonic initialization.  We design a progressive initialization strategy based on Harmonic bases of the volume mesh. By projecting the material vector into the Harmonic subspaces of different ranks, volumetric homogenization always finds a reliable initial value for the two-stage elasticity fitting procedure.
[…]
@@
For instance, to switch the optimization target to $\bm{q}$ in Eq. 21, we right multiply $\mathsf{diag}(\bm{H},\bm{H})$ to $\partial\bm{g}/\partial\bm{\gamma}$ and $(\partial\epsilon_{i}/\partial\bm{\gamma})^{\top}$, and sandwich $\bm{G}_{i}$ with $\mathsf{diag}(\bm{H},\bm{H})$ as
@@ 7.1. Projective Dynamics
The global step is a standard linear solve in the form of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{K}\bm{x}=\bm{b}}$ where we have
@@
The global step stands as the most expensive computation along the pipeline , as the Cholesky decomposition of $K$ may yield a large dense matrix that is not GPU-friendly.
@@ 7.2. Domain Decomposition
[…]
To accelerate the global step computation, we exploit the component mode synthesis (CMS) (Craig Jr and Hale, 1988; Craig Jr, 1985) to decompose the system into multiple domains (or components) and build a linear subspace at each domain to analyze the dynamic responses of complex structures. CMS was originally designed for linear structural analysis, and its generalization to nonlinear simulation remains an open research problem. Fortunately, the unique modality of PD allows us to apply CMS just at the global stage solve, so we can partition the mesh into multiple domains without aligning the decomposition with the variation of the underlying knit patterns. Unknown DOFs at each domain can be now grouped into internal DOFs and boundary DOFs. As the name suggests, the boundary DOFs interface with the neighbor domains while the internal DOFs are isolated by the boundary DOFs.
Figure 7. Comparison with HYLC (1$\times$1 rib).  We compare our method with HYLC and homogenize a square knitted fabric of the periodic 1$\times$1 rib pattern. The full-scale YLS results are shown at the bottom for reference. In general, both our method and HYLC yield plausible results, and both are visually similar to YLS results. Volumetric homogenization uses a mesh of $38$K DOFs, while HYLC only has $9$K DOFs. Nevertheless, our method takes 85ms per frame and is 180$\sim$ 180$\times$ 180 × faster than HYLC.
[…]
@@
Due to the linearity of this problem, $\bm{\Psi}_{(d)}^{(i)}$ encodes all the possible internal responses induced by boundary stimuli. When the domain does not undertake any non-boundary loads, $\bm{\Psi}_{(d)}^{(i)}$ relates boundary DOFs and internal DOFs as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{x}^{(i)}_{(d)}=\bm{\Psi}_{(d)}^{(i)}\bm{x}^{(b)}_{(d)}}$. They are a.k.a. boundary modes in CMS, and serve as subspace basis vectors for our global solve. For non-boundary responses, we compute a compact set of eigenvectors (e.g., 20) of $\bm{K}_{(d)}^{(ii)}$ corresponding to the smallest eigenvalues
@@
Figure 8. Comparison with HYLC (stockinette).  Given fabric with a periodic stockinette pattern, our method effectively captures spatially varying deformation leveraging material heterogeneity, particularly under significant stretch as highlighted. In the example, volumetric homogenization uses a mesh of $42$K DOFs, while HYLC has $12$K DOFs. Our method is 350$\sim$ 350$\times$ 350 × faster than HYLC.
[…]
Figure 9. Comparison with HYLC (draping).  We drape the pre-stretched ribbing patch on a sphere. Our method better replicates local deformation in the middle than HYLC. YLS simulation results are not used for elasticity fitting, showing our method’s generalizability.
[…]
Domain decomposition and CMS-like subspace construction techniques make our pipeline less sensitive to global matrix change e.g., due to material update or the presence of new constraints of collisions. After the reduced global solve is obtained, we use Eq. 34 to convert the generalized coordinate to the fullspace result. This result is then sent to a GPU-based full-space solver to make sure the global solution is accurate. In our implementation, we use the aggregated Jacobi method (A-Jacobi) proposed in (Lan et al., 2022b). A-Jacobi combines two or three Jacobi iterations into one iteration to fully exploit the parallel capacity of modern GPUs. As a result, mesh-level simulation is efficient is nearly interactive in many examples reported in this paper.
Note that a fast and numerically stable mesh simulation is critical to the entire pipeline–it is not only helpful for simulating the volumetric homogenized knitwear materials but also an indispensable module for elasticity fitting. Recall that our adjoint Gauss-Newton is only feasible when Eq. 14 is satisfied. This requirement suggests we need to find the equilibrium configuration of the mesh given the current material vector $\bm{\gamma}$. Whenever $\bm{\gamma}$ are updated during the elasticity fitting process, we have to re-assemble and factorize $\bm{x}^{M}_{i}$, which is the most expensive computation along the pipeline. Without a fast solver, elasticity fitting at such a high-dimension material space is infeasible.